OpenStudy (gamereternity):

What is the approximate area of the shaded region? A. 104.86 cm² B. 87.47 cm² C. 27.02 cm² D. 10.54 cm²

OpenStudy (gamereternity):

@Hero

OpenStudy (jtug6):

I'd recommend calculating the area of the circle and then subtracting it from the square. Do you know the area of a circle?

OpenStudy (gamereternity):

the area of the circle is 21.98

OpenStudy (jtug6):

Not quite. The area of a circle is pi * r^2. Our diamater is 7, so our radius is half that, which is 3.5. So what's pi(3.5^2)?

OpenStudy (gamereternity):

isnt the answer c

OpenStudy (jtug6):

Just work with me. The area of a circle is pi(3.5^2). What is that? Use pi = 3.14.

OpenStudy (gamereternity):

so that equals 21.98

OpenStudy (gamereternity):

7x7=49-21.98=c

OpenStudy (jtug6):

It's 38.48, not 21.98. pi(3.5^2) is pi(12.25) = 38.48 or so

OpenStudy (jtug6):

but your method of solving is correct. 49 is the area of the square so that minus 38.48 is roughly...?

OpenStudy (gamereternity):

3.5x3.14=10.99x2=21.98

OpenStudy (jtug6):

Youre taking the radius and multiplying it pi and then squaring that. It's actually pi times the radius squared.

OpenStudy (jtug6):